# Reciprocal Reward Influence Encourages Cooperation From Self-Interested Agents

### v1

@@ 4.2 Keeping Score with Influence Balances
The amount of influence that a Reciprocator is able to exert in a single time step is heavily environment-dependent, so that it may not always be possible to immediately reciprocate previous influences. To encourage reciprocation over extended timescales, we continuously accumulate a measure of net influence over sequences of actions. We draw inspiration from the “debit” mechanism $D_{i}$ used by approximate Markov tit-for-tat (Lerer & Peysakhovich, 2018, amTFT) to accumulate the advantage $A$ gained by an opponent $i$ compared to a cooperative counterfactual baseline, which they compute by recursively summing $A$ at every timestep.
@@ 4.3 Intrinsic Reciprocal Reward
If agent $i$ takes a series of actions that improves the expected return of the Reciprocator over the learned baseline expectation, i.e., produces a net positive influence balance, then the Reciprocator should be motivated to reinforce this behavior by exerting a reciprocal positive influence on agent $i$’s expected return $R_{i}$, in order to encourage a higher likelihood of that behavior during policy updates. Similar logic applies for detrimental deviations, negative influence balance, and subsequent reciprocal punishment. We then define the intrinsic reciprocal reward $r^{R}_{rc|i}(t)$ received by $rc$ as
@@
     Observe initial state $\boldsymbol{s_{0}}$
@@
        Observe $\boldsymbol{a}_{t},\boldsymbol{r_{t}},\boldsymbol{s}_{t+1}$
        Store transition tuple $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{M}$ and joint transition tuple $(\boldsymbol{s_{t}},\boldsymbol{a_{t}},\boldsymbol{r_{t}},\boldsymbol{s_{t+1}})$ in $\mathcal{H}$
@@ 4.4 Policy Optimization with Reciprocal Rewards
We use experience replay to periodically update target networks and iteratively update our counterfactual baseline estimates towards these target values (Mnih et al., 2015). This provides two key benefits: first, periodically updating these policy-dependent functions stabilizes training and allows us to approximately ignore their gradients, and therefore the gradient of the intrinsic reciprocal reward, with respect to the agents’ policy parameters (Wang et al., 2020). With this assumption, we are able to use to use standard policy gradient methods to train our agents to jointly optimize the combined extrinsic and intrinsic rewards.
@@
5 Experiments
We conduct experiments using two commonly-used SSDs of varied complexity to demonstrate the shaping abilities of Reciprocators against other types of learning agents. For IPD, we consider memory-1 iterated games as in Foerster et al. (2018) and Lu et al. (2022), following the proof from Press & Dyson (2012) that longer-memory strategies provide no advantage over strategies conditioned on shorter memories. We use two methods to evaluate head-to-head performance in IPD: allowing agents to directly differentiate through the analytic, closed-form solution of the game as originally derived in Foerster et al. (2018), and more standard batched policy rollouts for a fixed episode length. We append “analytic” and “rollout” to the game names to denote the evaluation method used. For Coins, we augment the observation with the time remaining in the episode to prevent state aliasing and stabilize learning from experience replay (Pardo et al., 2022).
5.1 Sequential Social Dilemmas
Iterated Prisoners’ Dilemma (IPD): The iterated prisoners’ dilemma (IPD) is a temporally-extended version of the classical thought experiment, in which two prisoners are given a choice to either stay silent/cooperate (C) or confess/defect (D) with rewards given in Table 1(a). The Pareto-optimal strategy is for both agents to cooperate by maintaining their silence, but the only Nash equilibrium (in the non-iterated, single-shot case) is mutual defection.
Coins: Coins is an $n$-player game introduced by Lerer & Peysakhovich (2018) as a temporally-extended version of the IPD. In this game, $n$ players move around in a fixed-size grid and collect colored coins. Agents receive a reward of 1 for collecting a coin, and are punished if another agent collects their coin such that, if agents collect coins indiscriminately where $P(\text{collect own coin})\approx 1/n$, the net expected reward is 0. We show example rewards for $n=2$ in Figure 1(b).
@@ 5.2 Baselines
Although we have discussed an extensive body of work related to cooperation in SSDs, we select these baselines because they focus specifically on the problem of shaping other agents’ policies during simultaneous learning.
Naïve Learner (NL): As previously defined, NLs optimize their expected return with respect only to their own policy parameters $\theta$. In this work, we implement NLs using standard policy gradient methods (Sutton et al., 1999), which perform updates of the form $\theta_{t+1}=\theta_{t}+\alpha\nabla_{\theta}J(\pi_{\theta})|_{\theta_{t}}$, where $\alpha$ is the learning rate and $\nabla_{\theta}J(\pi_{\theta})|_{\theta_{t}}$ is the gradient of the objective function with respect to the policy parameters $\theta_{t}$ at step $t$.
@@
Each entry is the average reward per episode achieved by the row agent against the column agent. We bold cooperative outcomes, which we define as both agents’ outcomes being better than -1.25, with a difference of no more than 0.10 between them. Standard error of the mean is less than 0.01 for all experiments. M-MAML results are averaged across 10 initial policies.
Figure 2: Representative run of a Reciprocator vs. an NL in IPD-Rollout. Reciprocal reward per step (left axis) and percentage of cooperation (staying silent) over the course of an episode (right axis).
@@
Model-Free Opponent Shaping (MFOS): As briefly discussed in the introduction, MFOS meta-learns over multiple episodes of policy updates in order to accomplish long-horizon opponent shaping. In Lu et al. (2022), MFOS is implemented with inner and outer policies, where the outer policy either directly outputs an inner policy to play in each episode or a conditioning vector which is elementwise multiplied with an inner policy vector, as done in IPD-Analytic and Coins, respectively.
@@ 5.3 Implementation Details
In IPD-Analytic, we differentiate directly through the analytic solution to the matrix game. For rollout-based experiments, we implement all policy gradient-based agents using actor-critic architectures trained with proximal policy optimization and a clipped surrogate objective (Schulman et al., 2017, PPO-Clip). Target networks to estimate the $Q$-values in Equation 1 are updated every $k$ episodes using uniformly sampled experience from a replay buffer (Lin, 1992). Additional hyperparameter values and network architecture details can be found in Appendix A.1. Shaded regions in plots depict one standard error of the mean.
@@ 6 Results
Here we summarize the results of our experiments, which include round-robin evaluations in IPD and demonstrations of Reciprocators’ ability to reach cooperation and shape NLs in Coins. Following work by Yu et al. (2022) showing that Meta-MAPG and LOLA-DICE agents fail to achieve significant results in Coins, even with a simplified shared reward, we do not use them as baselines in Coins.
@@ 6.1 IPD
IPD-Analytic: In the analytic form of IPD, we show that Reciprocators are able to reach cooperative equilibria with all other baselines despite being only a first-order method using vanilla gradient descent [Table 1]. Although the Reciprocator is extorted by MFOS to small extent (-1.06 vs. -0.98, respectively), we emphasize that MFOS relies on extensive liberties such as the ability to observe thousands of parallel training runs and roll out new meta-policies against opponents that are repeatedly reset back to randomly initialized policies. Additionally, as explained in Section 4, stochastic policy rollouts allow the Reciprocator to influence opponent returns differently for different action sequences and provide a stronger learning signal, motivating additional experiments using sampled rollouts.
IPD-Rollout: In IPD-Rollout, we show that Reciprocator is able to consistently shape an NL in IPD over the course of a single learning trajectory with limited rollout samples. We observe interesting learning dynamics wherein the Reciprocator’s policy repeatedly oscillates between lower and higher levels of cooperation. We show that these oscillations are driven by the opposing intrinsic and extrinsic rewards, and observe that the derivative of the rate of cooperation corresponds to the reciprocal reward [Figure 2]. Mechanistically, we see that the Reciprocator initially learns a cooperative tit-for-tat strategy guided by a strong reciprocal reward. However, as the NL’s policy becomes more deterministic and $a^{i}$ becomes predictable, the $VI^{\boldsymbol{\pi}}_{i|rc}$ component of the influence balance decreases. As the extrinsic reward begins to dominate, the Reciprocator essentially reverts into an NL, leading to exploitative defection. However, this exploitation of the NL agent $i$ causes the $VI^{\boldsymbol{\pi}}_{rc|i}$ term to become negative, leading to a negative reciprocal reward. This, combined with an increase in the NL’s frequency of defection (leading to an increase in $VI^{\boldsymbol{\pi}}_{i|rc}$), produces a reversal back to cooperative behavior that is then reinforced by positive reciprocal rewards.
Figure 3: Shaping an NL in Coins. Percentage of own coins collected by NL during training when facing each opponent (left) and coin counts across 32 steps for Reciprocator vs. NL (right).
@@ 6.2 Coins
In this temporally-extended social dilemma, we see that the Reciprocator is able to shape NL-PPO into picking up more of its own coins at rates equal to MFOS. This allows both agents to achieve a positive reward, while needing only a fraction of the samples to converge to this behavior (32,000 steps for Reciprocator vs. 512,000 steps for MFOS) [Figure 3]. We show that this change is driven by changes in coin preference rather than in total collection, with both agents collecting at near-optimal pace. This is in contrast to MFOS [Appendix B.2], which shapes the NL into collecting fewer coins.
When two Reciprocators are pitted against each other, we see that they are quickly able to learn a cooperative strategy of collecting their own coins [Figure 4], resulting in an average reward of $\sim 8$ per 32 steps without needing self-play. This significantly outperforms MFOS and the reported performance of LOLA with opponent modeling (LOLA-OM), which achieved an average reward of only $\sim 2$ per 32 steps according to Foerster et al. (2018). One concern is that the intrinsic reciprocal reward overpowers the extrinsic reward, effectively changing the motivational structure of the game into one that is fully-cooperative. To address this, we additionally recorded the cumulative reciprocal reward received in each episode over the course of training [Appendix B.1] to show that the reciprocal reward quickly decreases to a negliglible fraction of the total reward.
Figure 4: Head-to-head results in symmetric Coins (two agents of the same kind). Total reward per episode (left), percentage of own coins collected (right). Reciprocators and Naive Learner results are plotted on a scale of single episodes (bottom x-axis) whereas MFOS results are plotted on a scale of meta-episodes, where one meta-episode is equivalent to 16 episodes (top x-axis).
@@ 7 Limitations and Future Directions
Our current implementation of Reciprocators using naïve RL algorithms remains limited in that it seeks to maximize its compound return within single episodes rather than across multiple episodes. Using handpicked weights to balance intrinsic and extrinsic rewards opens the door for a suboptimal tradeoff between long-term opponent shaping and short-term return maximization. This is partially addressed by the counterfactual target baseline updates, which allow the Reciprocator to lower the magnitude of the reciprocal reward in response to opponent policies that remain stationary over multiple episodes [Figure 2]. Future work will focus on methods to evaluate reciprocation efficacy across multiple episodes and dynamically tune the balance between reciprocal and extrinsic rewards.
@@ 8 Conclusion and Broader Impacts
Emerging interest in cooperative AI has been led by approaches that endow agents with higher-order shaping capabilities. Although these agents exhibit cooperative behaviors when pitted against equivalent opponents, they readily manipulate and exploit agents with simpler learning rules or lower computational capabilities. This is a fundamentally undesirable outcome in partially-adversarial interactions between learning agents, with the potential to exacerbate existing computational resource gaps beyond out-of-the-box performance.
We presented Reciprocators, naïve learning agents equipped with an intrinsic reward for multi-agent settings that seeks to influence the behavior of other, simultaneously learning agents towards mutually beneficial outcomes. We showed that Reciprocators are able to shape the behavior of other agents in the environment over the course of training, and can produce prosocial behaviors in a variety of sequential social dilemmas. To the best of our knowledge, Reciprocators comprise the first class of learning algorithms to achieve cooperation during simultaneous learning between two independent agents without needing meta-learning methods, knowledge of other agents’ learning algorithms, or training schedules such as self-play or tracing procedures which control opponent selection and rely on the symmetric player roles. We believe that these results show a promising avenue forward for inducing cooperative outcomes from a diverse array of learning algorithms.
Appendix A Appendix / supplemental material
All experiments were run on an Nvidia 3070 GPU with 8 GB of VRAM. For IPD-Analytic and Coins experiments involving MFOS agents, we adapted the original code from Lu et al. (2022).
A.1 Hyperparameter Details
Each experiment was run 8 times, with a batch size of 8192 for IPD-Analytic and 2048 for IPD-Rollout and Coins.

### v2

@@ 4.2 Keeping Score with Influence Balances
The amount of influence that a Reciprocator is able to exert in a single time step is heavily environment-dependent, so that it may not always be possible to immediately reciprocate previous influences. To encourage reciprocation over extended timescales, we continuously accumulate a measure of net influence over sequences of actions. We draw inspiration from the “debit” tally used by approximate Markov tit-for-tat (Lerer & Peysakhovich, 2018, amTFT), which tracks the cumulative advantage gained by an opponent compared to a fully cooperative baseline policy known a priori.
@@ 4.3 Intrinsic Reciprocal Reward
If agent $i$ takes a series of actions that improves the expected return of the Reciprocator over a baseline estimate, i.e., produces a net positive influence balance, then the Reciprocator should be motivated to reinforce this behavior by exerting a reciprocal positive influence on agent $i$’s expected return $R_{i}$, in order to encourage a higher likelihood of that behavior during policy updates. Similar logic applies for detrimental deviations, negative influence balance, and subsequent reciprocal punishment. We then define the intrinsic reciprocal reward $r^{R}_{rc|i}(t)$ received by $rc$ as
@@
     Observe initial state $s_{0}$
@@
        Observe $\boldsymbol{a}_{t},\boldsymbol{r_{t}},s_{t+1}$
        Store transition tuple $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{M}$ and joint transition tuple $(s_{t},\boldsymbol{a_{t}},\boldsymbol{r_{t}},s_{t+1})$ in $\mathcal{H}$
@@ 4.4 Policy Optimization with Reciprocal Rewards
We use experience replay to periodically update target networks and iteratively update our counterfactual baseline estimates towards these target values (Mnih et al., 2015). This provides two key benefits: first, periodically updating these policy-dependent functions stabilizes training and allows us to approximately ignore their gradients, and therefore the gradient of the intrinsic reciprocal reward, with respect to the agents’ policy parameters (Wang et al., 2020). With this assumption, we are able to use standard policy gradient methods to train our agents to jointly optimize the combined extrinsic and intrinsic rewards.
@@
5 Experiments
We conduct experiments using two commonly used SSDs of varied complexity to demonstrate the shaping abilities of Reciprocators against other types of learning agents. For IPD, we consider memory-1 iterated games as in Foerster et al. (2018) and Lu et al. (2022), following the proof from Press & Dyson (2012) that longer-memory strategies provide no advantage over strategies conditioned on shorter memories. We use two methods to evaluate head-to-head performance in IPD: allowing agents to directly differentiate through the analytic, closed-form solution of the game as originally derived in Foerster et al. (2018), and more standard batched policy rollouts for a fixed episode length. We append “analytic” and “rollout” to the game names to denote the evaluation method used. For Coins, we augment the observation given to the critics and value influence estimators with the time remaining in the episode to prevent state aliasing and stabilize learning from experience replay (Pardo et al., 2022), but do not provide them as input to the policy networks.
5.1 Sequential Social Dilemmas
Iterated Prisoners’ Dilemma (IPD): The iterated prisoners’ dilemma (IPD) is a temporally extended version of the classical thought experiment, in which two prisoners are given a choice to either stay silent/cooperate (C) or confess/defect (D) with rewards given in Table 1(a). The Pareto-optimal strategy is for both agents to cooperate by maintaining their silence, but the only Nash equilibrium (in the non-iterated, single-shot case) is mutual defection.
Coins: Coins is a temporally extended variant of the IPD introduced by Lerer & Peysakhovich (2018). In this game, two coins spawn randomly in a fixed-size grid, with each coin corresponding to one of the two players (designated by color matching). Each player moves around the grid and receives a reward of 1 for collecting any coin, and a punishment of $-1$ if the other agent collects their coin. If agents collect coins indiscriminately where $P(\text{collect own coin})\approx 1/n$, the net expected reward is 0. We show example rewards for various scenarios in Figure 1(b).
@@ 5.2 Baselines
We select the following baselines because they focus specifically on the problem of shaping other agents’ policies during simultaneous learning and without modifications to the environment. In Coins, we exclude Meta-MAPG and LOLA-DICE as baselines following work by Yu et al. (2022) showing that neither method is able to achieve significant results, even with a simplified shared reward.
Naïve Learner (NL): As previously defined, NLs optimize their expected return with respect only to their own policy parameters $\theta$. In this work, we implement NLs using policy gradient methods (Sutton et al., 1999), which perform updates of the form $\theta_{t+1}=\theta_{t}+\alpha\nabla_{\theta}J(\pi_{\theta})|_{\theta_{t}}$, where $\alpha$ is the learning rate and $\nabla_{\theta}J(\pi_{\theta})|_{\theta_{t}}$ is the gradient of the objective function with respect to the policy parameters $\theta_{t}$ at step $t$.
@@
Each entry is the average reward per episode achieved by the row agent against the column agent. Standard error of the mean (SE) is less than 0.01 for all experiments and M-MAML results are averaged across the 10 initial policies provided by Lu et al. (2022).
Figure 2: Representative run of a Reciprocator vs. an NL in IPD-Rollout. Average reciprocal reward per step (left axis) and probability of cooperation (right axis) over the course of an episode.
@@
Model-Free Opponent Shaping (MFOS): As briefly discussed in the introduction, MFOS meta-learns over multiple episodes of policy updates in order to accomplish long-horizon opponent shaping. In Lu et al. (2022), MFOS is implemented with inner and outer policies, where the outer policy either directly outputs an inner policy to play in each episode or a conditioning vector which is element-wise multiplied with an inner policy vector, as done in IPD-Analytic and Coins, respectively.
@@ 5.3 Implementation Details
In IPD-Analytic, we differentiate directly through the analytic solution to the matrix game. For rollout-based experiments, we implement all policy gradient-based agents using actor-critic architectures trained with proximal policy optimization using a clipped surrogate objective (Schulman et al., 2017, PPO-Clip). Results including naïve learners trained with this algorithm are denoted by NL-PPO. Target networks to estimate the $Q$-values in Equation 1 are updated every $k$ episodes using uniformly sampled experience from a replay buffer (Lin, 1992). Additional hyperparameter values and network architecture details can be found in Appendix A.
Figure 3: Shaping an NL in Coins. Proportion of own coins collected by NL during training when facing each opponent (left) and coin counts by type for Reciprocator vs. NL (right). Reciprocator and NL-PPO results are plotted on a scale of single episodes (bottom axis) whereas MFOS results are plotted on a scale of meta-episodes, where one meta-episode contains 16 episodes (top axis).
@@ 6 Results
We evaluate the Reciprocators’ performance against a variety of baselines in a round-robin tournament for IPD-Analytic. In Coins, we assess opponent-shaping performance against NLs and against another agent of the same type, which we refer to as symmetric Coins. Shaded regions indicate standard error of the mean (SE) over eight random seeds. We do not perform full round-robin experiments in non-analytic environments for the following reasons: the Reciprocators’ complex and stochastic learning rule renders LOLA’s assumptions invalid and the extended time to convergence due to reciprocal behavior makes collecting meta-training episodes for MFOS computationally prohibitive.
@@ 6.1 IPD
IPD-Analytic: In the analytic form of IPD, we show that Reciprocators are able to reach cooperative equilibria with all other baselines, resisting exploitation by higher-order methods such as LOLA and MFOS despite being only a first-order method using vanilla gradient descent [Table 1]. Although the Reciprocator is extorted by MFOS to small extent (-1.06 vs. -0.98, respectively), we emphasize that MFOS relies on extensive liberties such as the ability to observe thousands of parallel training runs against their opponents and roll out pre-trained meta-policies against newly initialized agents.
IPD-Rollout: Stochastic policy rollouts allow the Reciprocator to influence opponent returns differently for different action sequences and provide a stronger learning signal, motivating additional experiments using sampled rollouts. In this setting, we show that Reciprocator is able to consistently shape an NL over the course of a single learning trajectory with limited rollout samples and observe interesting oscillations in the Reciprocator’s rate of cooperation. Figure 2 suggests that these oscillations are driven by the opposing intrinsic and extrinsic rewards, where the derivative of the rate of cooperation corresponds to the reciprocal reward value.
We provide the following explanation for this phenomenon: although the Reciprocator initially learns a cooperative tit-for-tat strategy guided by a strong reciprocal reward, as the NL’s policy becomes deterministic and $a^{i}$ becomes predictable, the $VI^{\boldsymbol{\pi}}_{i|rc}$ component of the influence balance decreases to 0. As the extrinsic reward begins to dominate, the Reciprocator essentially reverts into an NL, leading to exploitative defection. However, exploitation of the NL agent $i$ causes the $VI^{\boldsymbol{\pi}}_{rc|i}$ term, and subsequently the reciprocal reward, to become negative. Combined with an increase in the NL’s frequency of defection, which leads to an increase in $VI^{\boldsymbol{\pi}}_{i|rc}$, the intrinsic reward produces a reversal back to cooperative behavior that is reinforced by positive reciprocal rewards, and the cycle repeats.
Figure 4: Head-to-head results in symmetric Coins (two agents of the same kind). Total (extrinsic) reward per episode (left), proportion of own coins collected (right). Again, Reciprocator and NL-PPO results are plotted on a scale of episodes and MFOS results are plotted on a scale of meta-episodes.
@@ 6.2 Coins
In this temporally extended social dilemma, we see that the Reciprocator is able to shape NL-PPO into picking up more of its own coins at rates significantly higher than MFOS. This allows both agents to achieve a positive reward, while needing only a fraction of the samples to converge to mutually beneficial behavior (each meta-episode consists of 16 sequential episodes) [Figure 3]. We show that this change is driven by changes in coin preference rather than in total collection, with both agents collecting at near-optimal pace. This is in contrast to MFOS, which does not shape the NL towards cooperation, but rather uses its pretraining advantage to suppress opponent learning altogether (Khan et al., 2023) by collecting coins faster than the NL can reach them [Appendix B.2].
When two Reciprocators are pitted against each other, we see that they quickly learn a cooperative strategy of collecting their own coins [Figure 4], resulting in an average reward of $\sim 8$ per 32 steps without needing self-play. This significantly outperforms MFOS and the reported performance of LOLA with opponent modeling (LOLA-OM), which achieves an average reward of only $\sim 2$ per 32 steps according to Foerster et al. (2018). One concern is that the intrinsic reciprocal reward dominates the extrinsic reward, effectively changing the motivational structure of these SSDs into a cooperative game. To address this, we recorded the cumulative reciprocal reward received per episode [Appendix B.1] to show that it quickly decreases to a negligible fraction of the total reward, similar to the scale of the reciprocal reward in IPD in Figure 2.
@@ 7 Limitations and Future Directions
Our implementation of Reciprocators using naïve RL algorithms remains limited in that it seeks to maximize its compound return within single episodes rather than across multiple episodes. Using handpicked weights to balance intrinsic and extrinsic rewards opens the door for a suboptimal tradeoff between long-term opponent shaping and short-term return maximization. This is partially mitigated by the counterfactual target baseline updates, which allow the Reciprocator to lower the magnitude of the reciprocal reward in response to opponent policies that remain stationary over multiple episodes [Figure 2]. Future work will focus on methods to evaluate reciprocation efficacy across multiple episodes and dynamically tune the balance between reciprocal and extrinsic rewards.
In terms of evaluation, Khan et al. (2023) found that Coins does not require history to enable opponent shaping, since the current state is often indicative of past actions. Due to computational limitations, we were unable to assess our method’s performance on the suite of Spatio-Temporal Representations of Matrix Games (STORM) designed to test shaping over longer time horizons. We do note that Reciprocators capture both types of memory necessary to achieve shaping as identified by Khan et al. (2023): the counterfactual baselines trained on replay buffers serve as a way to capture inter-episode context, and the recurrent policies and influence balance capture intra-episode history. Therefore, our method can in theory generalize to STORM environments, although we leave this to future work.
@@ 8 Conclusion and Broader Impacts
Emerging interest in cooperative AI has been led by approaches that endow agents with higher-order shaping capabilities. Although these agents exhibit cooperative behaviors when pitted against equivalent opponents, they readily manipulate and exploit agents with simpler learning rules or lower computational capabilities. This is a fundamentally undesirable outcome in partially adversarial interactions between learning agents, with the potential to exacerbate existing computational resource gaps beyond out-of-the-box pretrained performance.
We presented Reciprocators, agents which seek to influence the behavior of other, simultaneously learning agents towards mutually beneficial outcomes. We showed that Reciprocators can both learn prosocial behaviors and induce them from other agents in a variety of sequential social dilemmas, while remaining resistant to exploitation by higher-order agents. To the best of our knowledge, Reciprocators represent the first class of reinforcement learning algorithms to achieve cooperation during simultaneous learning between two independent agents without needing meta-learning methods, knowledge of other agents’ learning algorithms, or pretraining routines such as self-play or tracing procedures to control opponent selection. We believe that these results show a promising avenue forward for inducing cooperative outcomes from a diverse array of learning opponents.
Acknowledgements
This work was supported by the following awards to JCK: National Institutes of Health DP2NS122037 and NSF CAREER 1943467.
Appendix A Experimental Details
All experiments were run on Nvidia 3070 GPUs with 8 GB of VRAM. Results were averaged over eight random seeds, with a batch size of 8192 for IPD-Analytic and 2048 for IPD-Rollout and Coins. For all experiments involving MFOS agents, we adapted the original code from Lu et al. (2022), leaving all architectural choices and hyperparameters the same.
The architecture for both the actor and the critic consists of a state encoder (two convolutional layers followed by a linear layer, with ReLU activations between layers), followed by hidden linear layers as detailed in the table below and a final output linear layer. The actor has a softmax activation to output a policy over the action space, whereas the critic simply outputs a scalar value estimate.